OpenStudy (anonymous):

Need help! Will fan and medal Segment AB has point A located at (6, 5). If the distance from A to B is 5 units, which of the following could be used to calculate the coordinates for point B? 5 = square root of the quantity of x minus 6 all squared plus y minus 5 all squared 5 = square root of the quantity of x minus 5 all squared plus y minus 6 all squared 5 = square root of the quantity of x plus 6 all squared plus y plus 5 all squared 5 = square root of the quantity of x plus 5 all squared plus y plus 6 all squared

OpenStudy (anonymous):

@Mehek14 @_greatmath7

OpenStudy (anonymous):

@babykel

OpenStudy (anonymous):

@Black_ninja123

OpenStudy (anonymous):

@babykel in the meantime would you like to help me out on another question?

OpenStudy (anonymous):

Depends on question sure

OpenStudy (anonymous):

Ok here goes

OpenStudy (anonymous):

Write an equation of a line that passes through the point (7, 3) and is parallel to the line y = negative 2 over 3x + 3. y = 3 over 2x − 3 y = 3 over 2x + 3 y = negative 2 over 3x + 23 over 3 y = negative 2 over 3x − 23 over 3

OpenStudy (anonymous):

I'm having so much trouble with this.

OpenStudy (anonymous):

holup

OpenStudy (anonymous):

@LynFran

OpenStudy (anonymous):

@DragonsGirl

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@Mehek14 @babykel

OpenStudy (anonymous):

ummmmmmm im not sure considering i dont know how to do this but i did this in my lesson and would go for b its make the most sense to me(for the second one)